OpenStudy (anonymous):

Chemical formula c3h4 has how many carbon-carbon double bonds are present in the molecule

OpenStudy (anonymous):

for this u use the degree of saturation formula. so 3 C=3x2 +2=8-4=4/2=2 double bonds